OpenStudy (anonymous):

Find the volume of a regular pentagonal prism with a 7in apothem, 12 inch side length and a 30in length, im really stuck on this so help would be awesome!<3

OpenStudy (anonymous):

here's a pic if it helps:)

OpenStudy (anonymous):

Volume of Prism = (5/2)ash = Ah where, a = apothem length, s = side, h = height, A=area

OpenStudy (anonymous):

oky so (5/2)ash should equal the area multiplied by the height?

OpenStudy (anonymous):

yup\[{5\over2}\times (apothem) \times side \times height = Area \times Height \]

OpenStudy (anonymous):

with side do you mean side length or the number of sides?

OpenStudy (anonymous):

annd would height be the diameter?

OpenStudy (anonymous):

7in apothem 30in length (if yu turn the prim up-side the length turns into height, right?) 12 inch side length = side

OpenStudy (anonymous):

oh okay so the length would be the height. annd why is it 5/2 at the beginning is that some sort of equation?

OpenStudy (anonymous):

hmm.. it's from deriving the equation... it's just like the area for a triagle = (1/2)bh

OpenStudy (anonymous):

oh is there an equation for a pentagonal prism?

OpenStudy (anonymous):

oh and in the equation you said it will = area* height, did you mean the area of the pentagon?

OpenStudy (anonymous):

OH! never mind i just didnt read ur answer right I get how you got 5/2 xD and this makes perfect sense now so thanks so much you're really smart!!

OpenStudy (anonymous):

haha thank's (:

OpenStudy (anonymous):

hey could you help me on a problem? i have it posted but i cant seem to get any responses. thanks<33

OpenStudy (anonymous):

sorry i was eating, i'll look for yur post and try to help

OpenStudy (anonymous):

Thanks so much C: and its the one with a triangular prism on top and a rectangular prism on bottom.

OpenStudy (anonymous):

you should re-post the question, i can't find it hehe

OpenStudy (anonymous):

haha okay I will thanks so much again:)

OpenStudy (anonymous):

did you find it? :)